# Deep Learning for Clouds and Cloud Shadow Segmentation in Methane Satellite and Airborne Imaging Spectroscopy

### v2

@@ Abstract
Effective cloud and cloud shadow detection is a critical prerequisite for accurate retrieval of concentrations of atmospheric methane (CH4) or other trace gases in hyperspectral remote sensing. This challenge is especially pertinent for MethaneSAT, a satellite mission launched in March 2024, to fill a significant data gap in terms of resolution, precision and swath between coarse-resolution global mappers and fine-scale point-source imagers of methane, and for its airborne companion mission, MethaneAIR. MethaneSAT delivers hyperspectral data at an intermediate spatial resolution ($\sim 100\times 400$  m), whereas MethaneAIR provides even finer resolution ($\sim 25$ m), enabling the development of highly detailed maps of concentrations that enable quantification of both the sources and rates of emissions. In this study, we use machine learning methods to address the cloud and cloud shadow detection problem for sensors with these high spatial resolutions. Cloud and cloud shadows in remote sensing data need to be effectively screened out as they bias methane retrievals in remote sensing imagery and impact the quantification of emissions. We deploy and evaluate conventional techniques—including Iterative Logistic Regression (ILR) and Multilayer Perceptron (MLP)—with advanced deep learning architectures, namely U-Net and a Spectral Channel Attention Network (SCAN) method. Our results show that conventional methods struggle with spatial coherence and boundary definition, affecting the detection of clouds and cloud shadows. Deep learning models substantially improve detection quality: U-Net performs best in preserving spatial structure, while SCAN excels at capturing fine boundary details. Notably, SCAN surpasses U-Net on MethaneSAT data, underscoring the benefits of incorporating spectral attention for satellite-specific features. Additionally, we combine the predictions of both U-Net and SCAN through a Convolutional Neural Network (CNN). This method achieves state-of-the-art performance on both MethaneAIR (78.50±3.08% F1) and MethaneSAT (78.80±1.28% F1) datasets with efficient inference (4.1 ms per 1,000 km2). This in-depth assessment of various disparate machine learning techniques, as applied to MethaneSAT and MethaneAIR imaging spectroscopic data at varying spatial resolutions, demonstrates the strengths and effectiveness of advanced deep learning architectures in providing robust, scalable solutions for clouds and cloud shadow screening towards enhancing methane emission quantification capacity of existing and next-generation hyperspectral missions.
@@
11footnotetext: Center for Data and AI, Universidad de Concepción, Edmundo Larenas 310, Concepción, Biobio 403000022footnotetext: AstroAI, Center for Astrophysics — Harvard & Smithsonian, 60 Garden st., Cambridge, MA 0213833footnotetext: Environmental Defense Fund, 257 Park Avenue South New York, New York 1001044footnotetext: Department of Earth and Planetary Sciences, Harvard University, 26 Oxford st., Cambridge, MA 0213855footnotetext: Department of Physics, Harvard University, 17 Oxford st., Cambridge, MA 02138
@@ I Introduction
Remote sensing has rapidly evolved as a key tool for quantifying emissions of critically-important greenhouse gases, including carbon dioxide (CO2) and methane (CH4). Among these, methane is especially critical: although it has a relatively short atmospheric lifetime of about twelve years, CH4 exhibits over 80 times the warming potential of CO2 during the first two decades after emission myhre_2013 ; Etminan_2016 . This makes methane an attractive target for near-term climate mitigation efforts Shindell_2012 .
@@
Methane mitigation efforts have advanced in the past few years with the Global Methane Pledge weforum2024methane signed by over 150 countries aiming to reduce anthropogenic methane emissions by 30% by year 2030. Sectorally, the Oil and Gas Decarbonization Charter (OGDC) has pledged to reduce methane emissions intensity down to 0.2% of production by 2030, which includes over 50 oil and gas companies accounting for more than 40% of global productionogdc_2023 . It is important to track the performance of critical methane mitigation targets, and remote sensing has emerged as an effective tool for globally monitoring methane emissions at scale.
@@
Two primary remote sensing strategies have emerged for monitoring methane: high-resolution imaging spectrometers, such as AVIRIS-NG, GHGSat, Carbon Mapper, and Sentinel-2, which excel at identifying point sources Frankenberg_2016 ; Varon_2021 ; Guanter_2021 , and low-resolution global mappers like TROPOMI and GOSAT, which offer daily or near-real-time regional coverage Veefkind_2012 ; Watine-Guiu_2023 . These hyperspectral missions have provided important data on global scale atmospheric methane concentrations and information on high-emitting point sources in targeted domains. However, there exists an observing and data gap at the scales of individual areas where emission quantification information is not readily available for assessing emissions at high spatial resolution, high precision and wider coverage - limiting our overall understanding of methane emission sources and magnitude.
@@
The MethaneSAT mission Rohrschneider , launched in March 2024, directly addresses this gap by combining fine-scale resolution, high-precision measurements with broad swath coverage of 220 km at nadir (extendable to over 400 km swath at greater off-nadir viewing geometries). The satellite features two imaging spectrometers: one targeting CH4 (1589–1686 nm) and another measuring O2 (1249–1305 nm), enabling precise retrieval of methane concentrations over oil and gas production basins and agricultural regions distributed globally. MethaneAIR, its airborne companion Conway_2024 , provides critical data for algorithm development and instrument validation, as well as ongoing mapping of methane emissions. Figure 1 illustrates the contrasting spatial coverage capabilities of MethaneSAT and MethaneAIR.
@@
Figure 1: Hyperspectral methane observations from MethaneSAT and MethaneAIR platforms showing typical mapping coverage and spatial extent. Left: MethaneSAT image at 1640.5 nm wavelength, captured on September 4 of 2024, covering approximately 220 km × 200 km area with 25 km scale bar. Right: MethaneAIR image at 1635.0 nm wavelength, captured on June 2 of 2023, covering approximately 100 km × 100 km area with 15 km scale bar.
@@
In this work, we investigate a range of machine learning algorithms for the detection of clouds and shadows in MethaneSAT and MethaneAIR hyperspectral imagery. We evaluate conventional models such as logistic regression Edisanter_2016 ; park_2023 ; Gautam_2023 and multilayer perceptrons (MLPs) BENEDIKTSSON_1993 ; Yang_1999 ; Tian_1999 ; Taravat_2015 ; gomezchova_2017 , as well as deep learning approaches including U-Net Tushar_2024 ; Ronneberger2015 ; Jiao_2020 ; Wieland_2019 ; Miao_2022 ; wright_2024 ; Li_2022 ; Tan_2023 and Spectral Channel Attention Network (SCAN), the latter designed to apply channel-wise attention mechanisms for hyperspectral band selection. Our results demonstrate that while conventional models offer computational efficiency, they struggle to resolve spatially complex or spectrally ambiguous regions. Deep networks, on the other hand, yield higher accuracy—U-Net excels in spatial consistency, while SCAN improves spectral boundary delineation, particularly for shadow detection.
To further improve performance, we develop ensemble architectures that fuse predictions from U-Net and SCAN. We explore both MLP- and CNN-based fusion strategies, with the latter achieving up to 8% improvements in F1 score over baseline models. These ensembles effectively combine spatial coherence and spectral attention to outperform individual models, especially in regions with complex topography or partial cloud cover.
Our findings reinforce the importance of robust cloud and shadow masking for hyperspectral CH4 retrieval and provide scalable solutions to enhance the fidelity of MethaneSAT and MethaneAIR products, that are aimed at advancing the quantification of emissions from individual oil/gas and agricultural regions around the world. More broadly, they offer potential applications and insights to other hyperspectral missions facing similar atmospheric challenges in remote sensing-based greenhouse gas monitoring.
@@ II Background
Hyperspectral semantic segmentation (HSS), also known as pixel-wise classification, aims to assign distinct labels to each pixel based on their spatial and spectral characteristics Ahmad_2021 ; Grewal_2022 ; Kumar_2024 . The rich spectral information contained in hyperspectral data, spanning multiple contiguous narrow bands, enables precise discrimination of materials and atmospheric features that may be indistinguishable in conventional imaging.
@@
Conventional approaches for HSS primarily relied on spectral-based classification, where each pixel is classified independently based on its spectral signature PAOLETTI2019279 ; BENEDIKTSSON_1993 ; Yang_1999 ; park_2023 . Methods like logistic regression Li_2010 ; park_2023 ; Gautam_2023 and multilayer perceptrons (MLPs) BENEDIKTSSON_1993 ; Yang_1999 ; Tian_1999 ; Taravat_2015 ; gomezchova_2017 operate directly on the spectral vectors, learning the complex relationships between spectral bands and target classes. While these spectral-only approaches can effectively leverage the rich spectral information, they fail to capture important spatial context and patterns that could significantly improve segmentation accuracy.
@@
To address this limitation, modern deep learning architectures have been developed to jointly exploit both spectral and spatial information PAOLETTI2019279 ; munishama_2015 ; Hu_2015 . The U-Net Ronneberger2015 ; Tushar_2024 architecture, originally designed for biomedical image segmentation Ronneberger2015 , has been adapted for multispectral cloud and shadow segmentation using Landsat 8 Jiao_2020 , Sentinel-2 Wieland_2019 ; Miao_2022 ; wright_2024 , MODIS Li_2022 , and Gaofen-1 Tan_2023 multisperspectral data. The architecture incorporates 2D convolutions that process both spatial neighborhoods and spectral bands simultaneously, while its encoder-decoder structure with skip connections enables multi-scale feature capture while preserving fine spatial details, making it particularly effective for semantic segmentation tasks.
@@
More recently, Vision Transformer-based architectures have emerged as a powerful alternative for HSS analysis Dosovitskiy_2020 . These models employ self-attention mechanisms to capture long-range dependencies in both spatial and spectral dimensions. Transformers have been also applied to cloud detection using Landsat 8 Li_2024 , Sentinel-2 Zhang_2023b ; Yan_2023 , MODIS Ge_2024 , and Gaofen-1 Tan_2023 multispectral images. Although Vision Transformers have shown promising results in various hyperspectral applications Hong_2021 , their computational complexity and large data requirements can be challenging for operational scenarios ahmad_2025 .
@@
Atmospheric monitoring of greenhouse gases using hyperspectral remote sensing has significantly advanced with missions such as AVIRIS-NG, EMIT, GHGSat, Carbon Mapper, TROPOMI, and MethaneSAT. These instruments detect and quantify greenhouse gases such as CH4 by leveraging spectral absorption features, primarily in the 1.6 micron and 2.3 micron methane absorption bands. Conventional hyperspectral imaging techniques, including matched filtering fuhrmann1992 , band-pass filtering and differential optical absorption spectroscopy (DOAS; Platt_2008 ), rely on spectral signatures to enhance detection (e.g. CH4 plumes against background variability Thompson_2015 ; Foote_2020 ; Ehret_2022 ; kumar_2023 ). Therefore, these methods often require expert-driven parameter tuning and manual inspection to reduce false positives Frankenberg_2016 ; ruzicka_2023 .
@@
Cloud and shadow segmentation in hyperspectral imagery presents a different set of challenges from CH4 detection, which benefits from well-characterized spectral absorption features. These atmospheric features exhibit high spectral variability across illumination conditions, surface reflectance, and sensor factors, complicating detection efforts zhu_2015 ; Wieland_2019 ; Chai_2019 . Cloud shadow detection is particularly difficult because dark surfaces like water bodies resemble shadows spectrally, shadowed areas contain diverse land cover types with broad spectral ranges, and thin clouds with high transmittance create shadows that mix with clear pixel characteristics zhai_2018 . Consequently, conventional spectral-based classification methods struggle to generalize across diverse scenes Wang_2024 . Influence of cloud shadows in the measured top-of-atmosphere radiances can lead to significant biases in atmospehric methane retrievals in turn affecting emission quantification, and therefore correcting for such artifacts is essential for accurately characterizing emission patterns.

### v3

@@ Abstract
Effective cloud and cloud shadow detection is a critical prerequisite for accurate retrieval of concentrations of atmospheric methane (CH4) or other trace gases in hyperspectral remote sensing. This challenge is especially pertinent for MethaneSAT, a satellite mission launched in March 2024, to fill a significant data gap in terms of resolution, precision and swath between coarse-resolution global mappers and fine-scale point-source imagers of methane, and for its airborne companion mission, MethaneAIR. MethaneSAT delivers hyperspectral data at an intermediate spatial resolution ($\sim 100\times 400$  m), whereas MethaneAIR provides even finer resolution ($\sim 25$ m), enabling the development of highly detailed maps of concentrations that enable quantification of both the sources and rates of emissions. In this study, we use machine learning methods to address the cloud and cloud shadow detection problem for sensors with these high spatial resolutions. Cloud and cloud shadows in remote sensing data need to be effectively screened out as they bias methane retrievals in remote sensing imagery and impact the quantification of emissions. We deploy and evaluate conventional techniques—including Iterative Logistic Regression (ILR) and Multilayer Perceptron (MLP)—with advanced deep learning architectures, namely U-Net and a Spectral Channel Attention Network (SCAN) method. Our results show that conventional methods struggle with spatial coherence and boundary definition, affecting the detection of clouds and cloud shadows. Deep learning models substantially improve detection quality: U-Net performs best in preserving spatial structure, while SCAN excels at capturing fine boundary details. Notably, SCAN surpasses U-Net on MethaneSAT data, underscoring the benefits of incorporating spectral attention for satellite-specific features. Additionally, we combine the predictions of both U-Net and SCAN through a Convolutional Neural Network (CNN). This ensemble method achieves the best performance on both MethaneAIR (F1: 78.50±3.08%) and MethaneSAT (F1: 78.80±1.28%) datasets, representing improvements of 2% and 10% over conventional methods (U-Net: 78.50±3.08% and 68.56±0.36% F1, respectively), while maintaining efficient inference (4.1 ms per 1,000 km²). This in-depth assessment of various disparate machine learning techniques, as applied to MethaneSAT and MethaneAIR imaging spectroscopic data at varying spatial resolutions, demonstrates the strengths and effectiveness of advanced deep learning architectures in providing robust, scalable solutions for clouds and cloud shadow screening towards enhancing methane emission quantification capacity of existing and next-generation hyperspectral missions.
@@
11footnotetext: Center for Data and AI, Universidad de Concepción, Edmundo Larenas 310, Concepción, Biobio 403000022footnotetext: Center for Astrophysics $|$ Harvard & Smithsonian, 60 Garden st., Cambridge, MA 0213833footnotetext: Environmental Defense Fund, 257 Park Avenue South New York, New York 1001044footnotetext: Department of Earth and Planetary Sciences, Harvard University, 26 Oxford st., Cambridge, MA 0213855footnotetext: Department of Physics, Harvard University, 17 Oxford st., Cambridge, MA 02138
@@ I Introduction
Remote sensing has rapidly evolved as a key tool for quantifying emissions of critically-important greenhouse gases, including carbon dioxide (CO2) and methane (CH4). Among these, methane is especially critical: although it has a relatively short atmospheric lifetime of about twelve years, CH4 exhibits over 80 times the warming potential of CO2 during the first two decades after emission Myhre et al. (2013); Etminan et al. (2016). This makes methane an attractive target for near-term climate mitigation efforts Shindell et al. (2012).
@@
Methane mitigation efforts have advanced in the past few years with the Global Methane Pledge World Economic Forum (2024) signed by over 150 countries aiming to reduce anthropogenic methane emissions by 30% by year 2030. Sectorally, the Oil and Gas Decarbonization Charter (OGDC) has pledged to reduce methane emissions intensity down to 0.2% of production by 2030, which includes over 50 oil and gas companies accounting for more than 40% of global productionCOP28 UAE (2023). It is important to track the performance of critical methane mitigation targets, and remote sensing has emerged as an effective tool for globally monitoring methane emissions at scale.
@@
Two primary remote sensing strategies have emerged for monitoring methane Chauhan and Raval (2025); Mehrdad and Du (2025): high-resolution imaging spectrometers, such as AVIRIS-NG, GHGSat, PRISMA, EMIT, EnMAP, Carbon Mapper, and Sentinel-2, which excel at identifying point sources Frankenberg et al. (2016); Varon et al. (2021); Guanter et al. (2021); Zhang et al. (2022); Roger et al. (2025); Xu et al. (2025); Smith et al. (2025); Bue et al. (2025); Mancoridis et al. (2025), and low-resolution global mappers like TROPOMI and GOSAT, which offer daily or near-real-time regional coverage Veefkind et al. (2012); Watine-Guiu et al. (2023). These hyperspectral missions have provided important data on global scale atmospheric methane concentrations and information on high-emitting point sources in targeted domains. However, there exists an observing and data gap at the scales of individual areas where emission quantification information is not readily available for assessing emissions at high spatial resolution, high precision and wider coverage – limiting our overall understanding of methane emission sources and magnitude.
@@
The MethaneSAT mission Rohrschneider et al. , launched in March 2024, directly addresses this gap by combining fine-scale resolution, high-precision measurements with broad swath coverage of 220 km at nadir (extendable to over 400 km swath at greater off-nadir viewing geometries). The satellite features two imaging spectrometers: one targeting CH4 (1589–1686 nm) and another measuring O2 (1249–1305 nm), enabling precise retrieval of methane concentrations over oil and gas production basins and agricultural regions distributed globally. MethaneAIR, its airborne companion Conway et al. (2024), provides critical data for algorithm development and instrument validation, as well as ongoing mapping of methane emissions. Figure 1 illustrates the contrasting spatial coverage capabilities of MethaneSAT and MethaneAIR.
@@
Figure 1: Hyperspectral methane observations from MethaneSAT and MethaneAIR platforms showing typical mapping coverage and spatial extent. Left: MethaneSAT image at 1640.5 nm wavelength, captured on September 4 of 2024, covering approximately 220 km × 200 km area. Right: MethaneAIR image at 1635.0 nm wavelength, captured on June 2 of 2023, covering approximately 100 km × 100 km area.
@@
Deep learning has driven substantial progress in cloud and shadow detection for satellite imagery, evolving from pixel-wise classifiers using multilayer perceptrons and random forests Gómez-Chova et al. (2017); Taravat et al. (2015) to fully convolutional architectures that exploit spatial context. U-Net Ronneberger et al. (2015) became a dominant backbone due to its encoder-decoder structure with skip connections, with Jeppesen et al. Jeppesen et al. (2019) demonstrating robust cloud masking across diverse Landsat scenes and Wieland et al. Wieland et al. (2019) showing strong generalization across Landsat-8 and Sentinel-2. More recently, attention mechanisms have been explored to capture spectral relationships in high-dimensional data Sun et al. (2020); Wright et al. (2024); Li et al. (2022), while semantic segmentation architectures have been widely adopted for related tasks including land cover mapping and change detection Ma et al. (2019); Chen et al. (2021, 2025). Despite these advances, most existing methods target broadband multispectral sensors, leaving their applicability to imaging spectrometers with hundreds of contiguous channels largely unexplored. Furthermore, cloud shadow detection remains comparatively understudied, as shadows exhibit subtler spectral signatures that are difficult to distinguish from dark surfaces.
In this work, we investigate a range of machine learning algorithms for the detection of clouds and shadows in MethaneSAT and MethaneAIR hyperspectral imagery. We evaluate conventional models such as logistic regression Lo and Ientilucci (2016); Park et al. (2023); Gautam and Nadda (2023) and multilayer perceptrons (MLPs) J. A. BENEDIKTSSON and ERSOY (1993); Yang (1999); Tian et al. (1999); Taravat et al. (2015); Gómez-Chova et al. (2017), as well as deep learning approaches including U-Net Tushar et al. (2024); Ronneberger et al. (2015); Jiao et al. (2020); Wieland et al. (2019); Miao et al. (2022); Wright et al. (2024); Li et al. (2022); Tan et al. (2023) and Spectral Channel Attention Network (SCAN), the latter designed to apply channel-wise attention mechanisms for hyperspectral band selection. Our results demonstrate that while conventional models offer computational efficiency, they struggle to resolve spatially complex or spectrally ambiguous regions. Deep networks, on the other hand, yield higher accuracy—U-Net excels in spatial consistency, while SCAN improves spectral boundary delineation, particularly for shadow detection.
To further improve performance, we develop ensemble architectures that fuse predictions from U-Net and SCAN. We explore both MLP- and CNN-based fusion strategies, with the latter achieving up to 10% improvements in F1 score over baseline models. These ensembles effectively combine spatial coherence and spectral attention to outperform individual models, especially in regions with complex topography or partial cloud cover.
In summary, this study presents the first benchmark of machine learning methods for cloud and shadow segmentation in MethaneSAT and MethaneAIR hyperspectral imagery. We develop SCAN (Spectral Channel Attention Network), a novel architecture employing channel-wise attention mechanisms for hyperspectral band selection, and a CNN-based ensemble that fuses spatial and spectral information to achieve the best overall performance. The motivation of this work is to enable operational efficiency with high accuracy: because our models operate directly on Level-1 radiance data, they can identify and filter cloud and shadow-contaminated pixels prior to Level-2 retrievals, bypassing the computationally expensive full retrieval pipeline for these pixels. This has significant implications for MethaneSAT’s mission to progress towards global methane reduction and is readily transferable to other hyperspectral platforms facing similar preprocessing bottlenecks. We release publicly annotated datasets for both MethaneSAT (262 scenes) and MethaneAIR (508 scenes), along with reproducible code for all evaluated models.
@@ II Background
Hyperspectral semantic segmentation (HSS), also known as pixel-wise classification, aims to assign distinct labels to each pixel based on their spatial and spectral characteristics Ahmad et al. (2021); Grewal et al. (2022); Kumar et al. (2024). The rich spectral information contained in hyperspectral data, spanning multiple contiguous narrow bands, enables precise discrimination of materials and atmospheric features that may be indistinguishable in conventional imaging.
@@
Conventional approaches for HSS primarily relied on spectral-based classification, where each pixel is classified independently based on its spectral signature Paoletti et al. (2019); J. A. BENEDIKTSSON and ERSOY (1993); Yang (1999); Park et al. (2023). Methods like logistic regression Li et al. (2010); Park et al. (2023); Gautam and Nadda (2023) and multilayer perceptrons (MLPs) J. A. BENEDIKTSSON and ERSOY (1993); Yang (1999); Tian et al. (1999); Taravat et al. (2015); Gómez-Chova et al. (2017) operate directly on the spectral vectors, learning the complex relationships between spectral bands and target classes. While these spectral-only approaches can effectively leverage the rich spectral information, they fail to capture important spatial context and patterns that could significantly improve segmentation accuracy.
@@
To address this limitation, modern deep learning architectures have been developed to jointly exploit both spectral and spatial information Paoletti et al. (2019); Munishamaiaha et al. (2023); Hu et al. (2015). The U-Net Ronneberger et al. (2015); Tushar et al. (2024) architecture, originally designed for biomedical image segmentation Ronneberger et al. (2015), has been adapted for multispectral cloud and shadow segmentation using Landsat 8 Jiao et al. (2020), Sentinel-2 Wieland et al. (2019); Miao et al. (2022); Wright et al. (2024), MODIS Li et al. (2022), and Gaofen-1 Tan et al. (2023) multisperspectral data. The architecture incorporates 2D convolutions that process both spatial neighborhoods and spectral bands simultaneously, while its encoder-decoder structure with skip connections enables multi-scale feature capture while preserving fine spatial details, making it particularly effective for semantic segmentation tasks.
@@
More recently, Vision Transformer-based architectures have emerged as a powerful alternative for HSS analysis Dosovitskiy et al. (2020). These models employ self-attention mechanisms to capture long-range dependencies in both spatial and spectral dimensions. Transformers have been also applied to cloud detection using Landsat 8 Li and Wang (2024), Sentinel-2 Zhang et al. (2023); Yan et al. (2023), MODIS Ge et al. (2024), and Gaofen-1 Tan et al. (2023) multispectral images. Although Vision Transformers have shown promising results in various hyperspectral applications Hong et al. (2021), their computational complexity and large data requirements can be challenging for operational scenarios Ahmad et al. (2025).
@@
Atmospheric monitoring of greenhouse gases using hyperspectral remote sensing has significantly advanced with missions such as AVIRIS-NG, EMIT, GHGSat, Carbon Mapper, TROPOMI, and MethaneSAT. These instruments detect and quantify greenhouse gases such as CH4 by leveraging spectral absorption features, primarily in the 1.6 micron and 2.3 micron methane absorption bands. Conventional hyperspectral imaging techniques, including matched filtering Fuhrmann et al. (1992), band-pass filtering and differential optical absorption spectroscopy (DOAS; Platt and Stutz (2008)), rely on spectral signatures to enhance detection (e.g. CH4 plumes against background variability Thompson et al. (2015); Foote et al. (2020); Ehret et al. (2022); Kumar et al. (2023)). Therefore, these methods often require expert-driven parameter tuning and manual inspection to reduce false positives Frankenberg et al. (2016); Růžička et al. (2023).
@@
Cloud and shadow segmentation in hyperspectral imagery presents a different set of challenges from CH4 detection, which benefits from well-characterized spectral absorption features. These atmospheric features exhibit high spectral variability across illumination conditions, surface reflectance, and sensor factors, complicating detection efforts Zhu et al. (2015); Wieland et al. (2019); Chai et al. (2019). Cloud shadow detection is particularly difficult because dark surfaces like water bodies resemble shadows spectrally, shadowed areas contain diverse land cover types with broad spectral ranges, and thin clouds with high transmittance create shadows that mix with clear pixel characteristics Zhai et al. (2018). Consequently, conventional spectral-based classification methods struggle to generalize across diverse scenes Wang et al. (2024). Influence of cloud shadows in the measured top-of-atmosphere radiances can lead to significant biases in atmospehric methane retrievals in turn affecting emission quantification, and therefore correcting for such artifacts is essential for accurately characterizing emission patterns.